if i have a regular hexagon with a perimeter of 216 inches, how many square meters of area will the hexagon be? please explain the steps . thank you.
Answer by rothauserc(4718)

The area of a hexagon can be calculated knowing the length of one of the sides,
area = (3/2) · sqrt(3) · s^2
now we know the perimeter is 216 inches, therefore a side is
s = 216 / 6 = 36 inches = 36 / 39.3700787 meters = 0.914400001 meters
note that there are 39.3700787 inches in a meter
area = (3/2) · sqrt(3) · (0.914400001)^2 = 2.172322608 square meters